4. The population of a certain village at one time was a perfect square. Later, with an increase of 100, the population was one more than a perfect square. Now, with an additional increase of 100, the population is again a perfect square. What was the original population? (Give all possible answers.)


1
Expert's answer
2021-03-02T05:12:00-0500

Let "n^2=" original population count. Then


"n^2+100=a^2+1"

"n^2+200=b^2"

"(a-n)(a+n)=99"

"(b-n)(b+n)=200"

"99=1\times99=3\times33=9\times 11"

"200=1\times200=2\times100=4\times50=5\times20"

"=8\times25=10\times20=8\times25"

"\begin{matrix}
 a-n=1 \\
 a+n=99
\end{matrix}=>\begin{matrix}
 n=49 \\
 a=50
\end{matrix}"

"49^2=2401,"

"49^2+100=2501=50^2+1,"

"49^2+200=2601=51^2"



"\begin{matrix}
 a-n=9 \\
 a+n=11
\end{matrix}=>\begin{matrix}
 n=15 \\
 a=18
\end{matrix}"

"15^2=225,"

"15^2+100=325=18^2+1,"

"15^2+200=425, 20^2=400, 21^2=441"

Does not satisfy.



"\begin{matrix}
 a-n=1 \\
 a+n=99
\end{matrix}=>\begin{matrix}
 n=1 \\
 a=10
\end{matrix}"

"1^2=1,"

"1^2+100=101=10^2+1,"

"1^2+200=201, 14^2=196, 15^2=225"

Does not satisfy.


The original population of a certain village was 2401 people.